# library.bib

### v2

@@ 2.9 Word maps
Write $A=1-(12+\epsilon)\alpha,$ which is $>0$ by our choice of $\epsilon.$ Then the RHS of (15) is equal to
$C(x,\delta)\left(\frac{C^{2}}{n^{A}}\right)\left(\frac{1-\left(\frac{C^{2}}{n^{A}}\right)^{n^{\alpha}-1}}{1-\left(\frac{C^{2}}{n^{A}}\right)}\right),$
@@
By [MAgeeDelaSalle, Lemma 5.14], there exist $y_{1},\dots,y_{m}\in\mathbb{C}\left[F_{r}\right]$ and $\delta^{\prime}=\delta^{\prime}(\delta),$ such that if, for every $i,$
@@ 4.1 Overview of proof
Rather than working directly with the character $\chi^{\lambda^{+}(n)},$ we use Corollary 2.11 so that instead our task is to compute the bitrace of $\mathcal{Q}_{\lambda,n}\circ w\left(g_{1},\dots,g_{r}\right)$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}.$
This amounts to counting the number of fixed points in $\left(\mathbb{C}^{n}\right)^{\otimes k}$ of this map. Recall that $\mathcal{Q}_{\lambda,n}^{2}=\mathcal{Q}_{\lambda,n}$ and that, since it is a linear combination of $p_{\pi},$ the action of $\mathcal{Q}_{\lambda,n}$ commutes with the action of any $g\in S_{n}$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}.$ With this in mind, if $w=x_{i_{1}}^{\pm 1}\dots x_{i_{l(w)}}^{\pm 1}$, what we are actually interested in is
$\mathrm{btr}_{\left(\mathbb{C}^{n}\right)^{\otimes k}}\left(g_{i_{1}}^{\pm 1}\circ\mathcal{Q}_{\lambda,n}\circ\dots\circ g_{i_{l(w)}}^{\pm 1}\circ\mathcal{Q}_{\lambda,n}\right),$
where each $g_{i}\in S_{n}$ is a permutation substituted in for the generator $x_{i}.$
@@
In §4.5, we prove a variant of a theorem of Louder and Wilton that relates the Euler characteristic of a graph with the number of immersed $w$–cycles (see [LouderWilton, Theorem 1.2]) to obtain our final bound for the expected character.
@@ 4.2 Expected character as a ratio of polynomials in $\frac{1}{n}$
In addition to our main theorem on word maps, we give another formulation for the expected stable irreducible character of a $w$–random permutation as the ratio of two polynomials in $\frac{1}{n}.$ This formulation is required for the methods in §3.
@@ 4.3 Combinatorial integration
Fix a word $w\in F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle.$ We will assume that $w$ is not the identity and that $w$ is cyclically reduced. We will also assume that every generator appears in $w$. If this were not the case, then we can relabel the $x_{i}$ that do appear in $w$ and write $w\in F_{m}=\left\langle x_{i_{1}},\dots,x_{i_{m}}\right\rangle$ for some $m<r$ and then apply the exact same proof.
@@
Suppose that $\{v_{p}\}$ is an orthonormal basis for $\mathcal{U}_{\lambda,n}\cong V^{\lambda^{+}(n)}\otimes V^{\lambda}.$ Then, for any $\left(g_{x_{1}},\dots,g_{x_{r}}\right)\in S_{n}^{r},$ we have
$\chi^{\lambda^{+}(n)}\left(w\left(g_{x_{1}},\dots,g_{x_{r}}\right)\right)=d_{\lambda}\mathrm{tr}_{\mathcal{U}_{\lambda,n}}\left(w\left(g_{x_{1}},\dots,g_{x_{r}}\right)\right),$
@@
$=$ $d_{\lambda}\sum_{p_{i}}\left\langle g_{f_{1}}^{\epsilon_{1}}v_{p_{2}},\ v_{p_{1}}\right\rangle\left\langle g_{f_{2}}^{\epsilon_{2}}v_{p_{3}},\ v_{p_{2}}\right\rangle\dots\left\langle g_{f_{l(w)}}^{\epsilon_{l(w)}}v_{p_{1}},\ v_{p_{l(w)}}\right\rangle.$
@@
Recall that $\mathcal{U}_{\lambda,n}\subset D_{k}(n),$ so that we may assume that the above sum is over all $I=\left(i_{1},\dots,i_{k}\right)$ with all indices distinct. This is an important observation.
@@
In our expression, we will simply write $(I/J)$ in place of $(I/J)(\epsilon)$ where it is clear which epsilon we are considering. To be even more clear, in the sum below, for each $f_{i},$ if $\epsilon_{i}=1,$ then the corresponding inner product has $e_{I_{f_{i}}^{z}}$ on the LHS, where $z\in\{1,\dots,|w|_{f_{i}}\}$ denotes the number of times $f_{i}$ has appeared in the subword $f_{1}^{\epsilon_{1}}\dots f_{i}^{\epsilon_{i}}.$ The corresponding pair of $\beta$–terms has $\beta_{p,I},$ with $\beta_{p,J}$ conjugated. If $\epsilon_{i}=-1,$ we swap the positions of $I$ and $J$ [^3] We use this notation so that, in the graph construction detailed in Section 4.4, $\left(J_{f}^{i}\right)_{j}$ always represents the initial vertex of an $f$–edge and $\left(I_{f}^{i}\right)_{j}$ always represents the terminal vertex.
@@
$d_{\lambda}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)\left\langle g_{f_{1}}^{\epsilon_{1}}e_{\left(I_{f_{1}}^{1}/J_{f_{1}}^{1}\right)},\ e_{\left(J_{f_{1}}^{1}/I_{f_{1}}^{1}\right)}\right\rangle\dots$ (20)
@@
$d_{\lambda}\sum_{p_{i}}\sum_{\begin{aligned} I_{x_{1}}^{1},I_{x_{1}}^{2},I_{x_{2}}^{1},I_{x_{2}}^{2},\\ J_{x_{1}}^{1},J_{x_{1}}^{2},J_{x_{2}}^{1},J_{x_{2}}^{2}\end{aligned}}\left(\beta_{p_{2},I_{x_{1}}^{1}}\right)\left(\bar{\beta}_{p_{1},J_{x_{1}}^{1}}\right)\left(\beta_{p_{3},I_{x_{2}}^{1}}\right)\left(\bar{\beta}_{p_{2},J_{x_{2}}^{1}}\right)\left(\beta_{p_{4},J_{x_{1}}^{2}}\right)\left(\bar{\beta}_{p_{3},I_{x_{1}}^{2}}\right)$
@@
$d_{\lambda}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)\prod_{f\in\{x_{1},\dots,x_{r}\}}\int_{g_{I^{f},J^{f}}},$ (21)
@@
For each $f\in\{x_{1},\dots,x_{r}\}$ and for each fixed collection of multi–indices $I_{f}^{1},\dots,I_{f}^{|w|_{f}},J_{f}^{1},\dots,J_{f}^{|w|_{f}}$, this integral can be computed using the Weingarten calculus for the symmetric group. Using Theorem 2.14 with $m=|w|_{f}k,$ we have
@@
To each partition we associate a diagram (in a similar way to §2.2) – the diagram has $|w|_{f}$ rows, each containing exactly $k$ vertices. The vertices in row $R$ are labeled from $(R-1)k+1$ to $Rk$, and two vertices are connected if and only if the corresponding vertex labels are in the same block of of the partition.
We will also refer to the diagrams corresponding to $\sigma_{f}$ and $\tau_{f}$ as $\sigma_{f}$ and $\tau_{f}$ where this does not cause confusion.
@@
The next lemma asserts that every vertex must be connected to at least one other vertex
@@
$d_{\lambda}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)$
@@
This is exactly the conjugate of the coefficient of $e_{(J_{f}^{1})_{1}}\otimes\dots\otimes e_{(J_{f}^{1})_{q-1}}\otimes e_{(J_{f}^{1})_{q+1}}\otimes\dots e_{(J_{f}^{1})_{k}}$ in $T_{q}\left(v_{p_{1}}\right).$ But then, $v_{p_{1}}$ belongs to an orthonormal basis for $\mathcal{U}_{\lambda,n},$ and $\mathcal{U}_{\lambda,n}\subset A_{k}(n).$ In particular, $v_{p}\in\mathrm{Ker}\left(T_{q}\right),$ so that
@@
Remark 4.5.
It follows that if any of $x_{1},\dots,x_{r}$ appear only once in $w$, then there are no such set partitions in $\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right)$, and so in these cases $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=0.$ Such words are primitive and we know from [HananyPuder] that the expected stable irreducible character for primitive words is zero, unless $k=0.$ From now on, we will only be considering non–primitive words.
@@
$\sigma_{f}\leftrightarrow J_{f}^{\Sigma}$
@@
$d_{\lambda}\sum_{\sigma_{f},\tau_{f}}^{\star}\left(\prod_{f\in\{x_{1},\dots,x_{r}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)\sum_{p_{1}}\sum_{\begin{aligned} \sigma_{f}\leftrightarrow J_{f}^{\Sigma},\\ \tau_{f}\leftrightarrow I_{f}^{\Sigma}\end{aligned}}\left(\prod_{f,w}\beta\right).$ (22)
This expression can be further simplified using the projection $\mathcal{Q}_{\lambda,n}.$ Consider Example 4.2. For all terms fixed in the expression except for, say, $p_{2},$ the sum that follows is
@@
$=$ $\sum_{\pi\in\mathrm{Part}\left([2k]\right)}c(n,k,\lambda,\pi)\left\langle p_{\pi}\left(e_{I_{x_{1}}^{1}}\right),e_{J_{x_{2}}^{1}}\right\rangle.$
@@
$d_{\lambda}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{w}}\left(\prod_{i=1}^{4}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},x_{2}\}}\mathrm{Wg}_{n,(|w|_{f}k)}\left(\sigma_{f},\tau_{f}\right)\right)$ (23)
@@
Now, recall that the multi–indices in our sum must have all indices distinct. Therefore, if $\pi_{1}$ has any two vertices in the top row connected, or any two vertices in the bottom row connected, then the term
$\left\langle p_{\pi_{1}}\left(e_{I_{x_{1}}^{1}}\right),e_{J_{x_{2}}^{1}}\right\rangle=0$
for any pair of multi–indices. Similarly for $\pi_{2},\ \pi_{3}$ and $\pi_{4}.$ So the only summands that contribute to (23) are those in which each $\pi_{i}\leq\hat{\pi}_{i},$ where $\hat{\pi}_{i}=\iota(\rho)$ for some permutation $\rho\in S_{k}.$ In this case, we can replace the sum over the inner product terms in (23) by the number of multi–indices $I_{x_{1}}^{1},I_{x_{1}}^{2},I_{x_{2}}^{1},I_{x_{2}}^{2},J_{x_{1}}^{1},J_{x_{1}}^{2},J_{x_{2}}^{1},J_{x_{2}}^{2}$ that simultaneously satisfy $\sigma_{f}\leftrightarrow J_{f}^{\Sigma},\ \tau_{f}\leftrightarrow I_{f}^{\Sigma}$ as well as all of the inner product terms being equal to one.
We write
$\mathrm{Part}_{\leq S_{k}}\left([2k]\right)\overset{\mathrm{def}}{=}\big{\{}\pi\in\mathrm{Part}\left([2k]\right):\ \pi\leq\iota(\rho),\ \rho\in S_{k}\big{\}}.$
Additionally, we will write
$\sum_{\pi}^{\leq S_{k}}$
to indicate that we are summing over $\pi\in\mathrm{Part}_{\leq S_{k}}\left([2k]\right).$
This argument applies for any non–identity, cyclically reduced, non–primitive word $w.$ With this notation and the definition below, this proves Theorem 4.7.
@@
$=$ $d_{\lambda}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},\dots,x_{r}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)$
$\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right).$
Combining (24) with Remark 2.12 and Lemma 2.15, the following bound is immediate.

### v3

@@ 2.9 Word maps
Write $A=1-(20+\epsilon)\alpha,$ which is >0>0 by our choice of $\epsilon.$ Then the RHS of (14) is equal to
$C(z,\delta)\left(\frac{C^{2}}{n^{A}}\right)\left(\frac{1-\left(\frac{C^{2}}{n^{A}}\right)^{n^{\alpha}-1}}{1-\left(\frac{C^{2}}{n^{A}}\right)}\right),$
@@
By [MdlS24, Lemma 5.14], there exist $y_{1},\dots,y_{m}\in\mathbb{C}\left[F_{r}\right]$ and $\delta^{\prime}=\delta^{\prime}(\delta),$ such that if, for every $i,$
@@ 4.1 Overview of proof
Rather than working directly with the character $\chi^{\lambda^{+}(n)},$ we use Corollary 2.11 so that instead our task is to compute the trace of $\mathcal{Q}_{\lambda,n}\circ w\left(g_{1},\dots,g_{r}\right)$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}.$
This amounts to counting the number of standard basis vectors of $\left(\mathbb{C}^{n}\right)^{\otimes k}$ fixed by this map. Recall that $\mathcal{Q}_{\lambda,n}^{2}=\mathcal{Q}_{\lambda,n}$ and that, since it is a linear combination of $p_{\pi},$ the action of $\mathcal{Q}_{\lambda,n}$ commutes with the action of any $g\in S_{n}$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}.$ With this in mind, if $w=x_{1}x_{2}x_{1}^{-1}x_{2}^{-1}\in F_{2}$ for example, what we are actually interested in is
$\mathrm{tr}_{\left(\mathbb{C}^{n}\right)^{\otimes k}}\left(g_{1}\circ\mathcal{Q}_{\lambda,n}\circ g_{2}\circ\mathcal{Q}_{\lambda,n}\circ g_{1}^{-1}\circ\mathcal{Q}_{\lambda,n}\circ g_{2}^{-1}\circ\mathcal{Q}_{\lambda,n}\right),$
where each $g_{i}\in S_{n}$.
@@
In §4.5, we prove a variant of a theorem of Louder and Wilton that relates the Euler characteristic of a graph with the number of immersed ww–cycles (see [LW17, Theorem 1.2]) to obtain our final bound for the expected character.
@@ 4.2 Expected character as a ratio of polynomials in $\frac{1}{n}$
In addition to our main theorem on word maps, we give another formulation for the expected stable irreducible character of a ww–random permutation as the ratio of two polynomials in $\frac{1}{n}.$ This formulation is required for the methods in §3.
@@ 4.3 Combinatorial integration
Fix a word $w\in F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle.$ We will assume that $w$ is not the identity and is not primitive and that $w$ is cyclically reduced. We will also assume that every $x_{i}$ appears at least twice in $w$. If $x_{j}$ did not appear in $w,$ then we could consider $w\in F_{r-1}=\left\langle x_{1},\dots,x_{j-1},x_{j+1},\dots,x_{r}\right\rangle$ and then proceed. If any $x_{j}$ appeared only once, then $w$ would be primitive and so $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=0.$
@@
Suppose that $\{v_{p}\}$ is an orthonormal basis for $\mathcal{U}_{\lambda^{+}(n)}\cong V^{\lambda^{+}(n)}\otimes V^{\lambda}.$ Then, for any $\left(g_{x_{1}},\dots,g_{x_{r}}\right)\in S_{n}^{r},$ we have
$\chi^{\lambda^{+}(n)}\left(w\left(g_{x_{1}},\dots,g_{x_{r}}\right)\right)=\frac{1}{d_{\lambda}}\mathrm{tr}_{\mathcal{U}_{\lambda^{+}(n)}}\left(w\left(g_{x_{1}},\dots,g_{x_{r}}\right)\right),$
@@
$=$ $\frac{1}{d_{\lambda}}\sum_{p_{i}}\left\langle g_{f_{1}}^{\epsilon_{1}}v_{p_{2}},\ v_{p_{1}}\right\rangle\left\langle g_{f_{2}}^{\epsilon_{2}}v_{p_{3}},\ v_{p_{2}}\right\rangle\dots\left\langle g_{f_{l(w)}}^{\epsilon_{l(w)}}v_{p_{1}},\ v_{p_{l(w)}}\right\rangle.$
@@
Recall that $\mathcal{U}_{\lambda^{+}(n)}\subset D_{k}(n),$ so that we may assume that the above sum is over all $I=\left(i_{1},\dots,i_{k}\right)$ with all indices distinct. This is an important observation.
@@
In our expression, we will simply write $(I/J)$ in place of $(I/J)(\epsilon)$ where it is clear which epsilon we are considering. To be even more clear, in the sum below, for each $f_{i},$ if $\epsilon_{i}=1,$ then the corresponding inner product has $e_{I_{f_{i}}^{z}}$ on the LHS, where $z\in\{1,\dots,|w|_{f_{i}}\}$ denotes the number of times $f_{i}$ has appeared in the subword $f_{1}^{\epsilon_{1}}\dots f_{i}^{\epsilon_{i}}.$ The corresponding pair of $\beta$–terms has $\beta_{p,I},$ with $\beta_{p,J}$ conjugated. If $\epsilon_{i}=-1,$ we swap the positions of $I$ and JJ. [^4] We use this notation so that, in the graph construction detailed in Section 4.4, $\left(J_{f}^{i}\right)_{j}$ always represents the initial vertex of an ff–edge and $\left(I_{f}^{i}\right)_{j}$ always represents the terminal vertex.
@@
$\frac{1}{d_{\lambda}}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)\left\langle g_{f_{1}}^{\epsilon_{1}}e_{\left(I_{f_{1}}^{1}/J_{f_{1}}^{1}\right)},\ e_{\left(J_{f_{1}}^{1}/I_{f_{1}}^{1}\right)}\right\rangle\dots$ (19)
@@
$\frac{1}{d_{\lambda}}\sum_{p_{i}}\sum_{\begin{aligned} I_{x_{1}}^{1},I_{x_{1}}^{2},I_{x_{2}}^{1},I_{x_{2}}^{2},\\ J_{x_{1}}^{1},J_{x_{1}}^{2},J_{x_{2}}^{1},J_{x_{2}}^{2}\end{aligned}}\left(\beta_{p_{2},I_{x_{1}}^{1}}\right)\left(\bar{\beta}_{p_{1},J_{x_{1}}^{1}}\right)\left(\beta_{p_{3},I_{x_{2}}^{1}}\right)\left(\bar{\beta}_{p_{2},J_{x_{2}}^{1}}\right)\left(\beta_{p_{4},J_{x_{1}}^{2}}\right)\left(\bar{\beta}_{p_{3},I_{x_{1}}^{2}}\right)$
@@
$\frac{1}{d_{\lambda}}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)\prod_{f\in\{x_{1},\dots,x_{r}\}}\int_{g_{I^{f},J^{f}}},$ (20)
@@
For each $f\in\{x_{1},\dots,x_{r}\}$ and for each fixed collection of multi–indices $I_{f}^{1},\dots,I_{f}^{|w|_{f}},J_{f}^{1},\dots,J_{f}^{|w|_{f}}$, this integral can be computed using the Weingarten calculus for the symmetric group. Using (2.14) we have
@@
To each partition we associate a diagram (in a similar way to §2.2) – the diagram has $|w|_{f}$ rows, each containing exactly $k$ vertices. The vertices in row $R$ are labeled from (R1)k+1(R-1)k+1 to $Rk$, and two vertices are connected if and only if the corresponding vertex labels are in the same block of of the partition. We will also refer to the diagrams corresponding to $\sigma_{f}$ and $\tau_{f}$ as $\sigma_{f}$ and $\tau_{f}$ where this does not cause confusion.
@@
The next lemma asserts that every vertex of $\sigma_{f}$ and $\tau_{f}$ must be connected to at least one other vertex.
@@
$\frac{1}{d_{\lambda}}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)$
@@
This is exactly the conjugate of the coefficient of $e_{(J_{f}^{1})_{1}}\otimes\dots\otimes e_{(J_{f}^{1})_{q-1}}\otimes e_{(J_{f}^{1})_{q+1}}\otimes\dots e_{(J_{f}^{1})_{k}}$ in $T_{q}\left(v_{p_{1}}\right).$ But then, $v_{p_{1}}$ belongs to an orthonormal basis for $\mathcal{U}_{\lambda,n},$ and $\mathcal{U}_{\lambda,n}\subset A_{k}(n).$ In particular, $v_{p}\in\ker\left(T_{q}\right),$ so that
@@
$J_{f}^{\Sigma}\leftrightarrow\sigma_{f}$
@@
$\frac{1}{d_{\lambda}}\sum_{\sigma_{f},\tau_{f}}^{\star}\left(\prod_{f\in\{x_{1},\dots,x_{r}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)\sum_{p_{i}}\sum_{\begin{aligned} J_{f}^{\Sigma}\leftrightarrow\sigma_{f}\\ I_{f}^{\Sigma}\leftrightarrow\tau_{f}\end{aligned}}\left(\prod_{f,w}\beta\right).$ (21)
This expression can be further simplified using the projection $\mathcal{Q}_{\lambda,n}.$ Consider Example 4.2 – for all terms fixed in the expression except for e.g. $p_{2},$ the ensuing sum is
@@
$=$ $\sum_{\pi}^{\leq S_{k}}c(n,k,\lambda,\pi)\left\langle p_{\pi}\left(e_{I_{x_{1}}^{1}}\right),e_{J_{x_{2}}^{1}}\right\rangle.$
@@
$\frac{1}{d_{\lambda}}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{i=1}^{4}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},x_{2}\}}\mathrm{Wg}_{n,(|w|_{f}k)}\left(\sigma_{f},\tau_{f}\right)\right)$ (22)
@@
The same argument applies for any non–identity, non–primitive, cyclically reduced word $w\in F_{r}.$
@@
With this notation, we have proved the following theorem.
@@
$=$ $\frac{1}{d_{\lambda}}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},\dots,x_{r}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right).$
Combining (23) with Lemma 2.12 and Remark 2.15, the following bound is immediate.